During a football game, a receiver has just caught a pass and is standing still. Before he can move, a tackler, running at a velocity of +3.1 m/s, grabs and holds onto him so that they move off together with a velocity of +1.7 m/s. If the mass of the tackler is 120 kg, determine the mass of the receiver. Assume momentum is conserved.
kg
Answer #1

Let mass of receiver = m

From conservation of momentum,

m1v1 = (m1 + m)*v2

120*3.1 = (120 + m)*1.7

m = 98.82 kg